OpenStudy (anonymous):

Algebra Help: 3 plus the reciprocal of a number equals 7 divided by that number. What is the number?

OpenStudy (anonymous):

3 + 1/n = 7/n 3 = 7/n - 1/n 3 = 6/n n = 2

OpenStudy (anonymous):

wait does / mean divide or r u seperatin the parts of the question

OpenStudy (anonymous):

divide (sorry for being confusing)

OpenStudy (anonymous):

no its ok and thnx :)